OpenStudy (babyslapmafro):

Please help me integrate the following problem: sin(4x)sin(8x)dx

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\sin(4x)\sin(8x)dx\]

myininaya (myininaya):

Ok so let's recall a couple of identities that include something like what you have inside that integral The cos sum and difference identity has something that looks like that. \[\cos(4x+8x)=\cos(4x)\cos(8x)-\sin(4x)\sin(8x)\] \[\cos(4x-8x)=\cos(4x)\cos(8x)+\sin(4x)\sin(8x)\] now we can solve this system of sin(4x)sin(8x) but multiplying the first identity by negative 1 and then adding the equals together like so: \[-\cos(4x+8x)=-\cos(4x)\cos(8x)+\sin(4x)\sin(8x)\] \[\cos(4x-8x)=\cos(4x)\cos(8x)+\sin(4x)\sin(8x)\] now add together \[-\cos(4x+8x)+\cos(4x-8x)=2\sin(4x)\sin(8x)\] or just \[\sin(4x)\sin(8x)=\frac{1}{2}(-\cos(12x)+\cos(-4x))\]

myininaya (myininaya):

There is actually an identity for this but I'm just showing you the identity using other identities. :)